In 1995, the life expectancy of males in a certain country was 66.9 years. In 2002, it was 69.7 years. Let E represent the life expectancy in year t and let t represent the number of years since 1995.
Find a linear function E(t) that fits the data.
E(t)= ?t+?
HELP PLEASE!!!

Answer by stanbon(75887)

In 1995, the life expectancy of males in a certain country was 66.9 years. In 2002, it was 69.7 years. Let E represent the life expectancy in year t and let t represent the number of years since 1995.
Find a linear function E(t) that fits the data.
---------------------------------
You have two points: (0,66.9) and (7,69.7)
slope = (69.7-66.9)/7 = 2/5
---------------------------
Form E = mt+b
You have y,x,and m; solve for "b":
66.9 = (2/5)(0)+b
b = 66.9
EQUATION:
E(t) = (2/5)t +66.9
Cheers,
Stan H.